Circular No. 8075

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003au IN NGC 6095 Further to IAUC 8070, M. Schwartz and W. Li report the LOTOSS

discovery of an apparent supernova (mag ∼ 19.2) on an unfiltered images taken on Feb. 17.5 and 18.5 UT. SN 2003au is located at α = 16^{h}11^{m}11^{s}.56, δ = +61^{o}15′58′′.2 (equinox 2000.0), which is 4′′.8 east and 6′′.9 south of the nucleus of NGC 6095. A KAIT image taken on 2002 June 5.3 showed nothing at this position (limiting mag ∼ 20.0).

S/2003 (1509) 1 W. J. Merline, Southwest Research Institute (SRI); L. M. Close, Uni-

versity of Arizona; P. M. Tamblyn, Binary Astronomy and SRI; F. Menard, Observatoire de Grenoble; C. R. Chapman, SRI; C. Dumas, Jet Propulsion Laboratory (JPL); G. Duvert, Observatoire de Grenoble; W. M. Owen, JPL; D. C. Slater, SRI; and M. F. Sterzik, European Southern Observatory (ESO), report the discovery on Feb. 13.3 UT of a satellite of (1509) Esclangona (V = 15.0) from K_{s}-band direct imaging with the ESO 8-m Very Large Telescope UT4/YEPUN (+ NAOS/CONICA adaptive optics system) on Cerro Paranal. The satellite was observed on 5 epochs spanning 3 nights. On Feb. 13.3214, the companion was at separation 0′′.20 (projected separation 140 km) in p.a. 292^{o}. The K_{s} brightness difference is ∼ 2.4 mag, giving an estimated diameter of the satellite of ∼ 4 km.

M87 E. Perlman, University of Maryland, Baltimore County; and J. Biretta

and W. Sparks, Space Telescope Science Institute, report strong flaring in a knot in the M87 jet: “Knot HST-1 is located 0′′.8 from the nucleus of M87 along p.a. −69^{o}.5. Flaring activity has been seen in our Hubble Space Telescope (HST) monitoring, which shows a steady, monotonic increase of the optical flux from 2001 August through 2002 Dec. 22, when the 220-nm flux peaked at 98 μJy, comparable to the nucleus in that band. These fluxes translate to U = 16.0 if a power-law spectrum (ν^{−0.6}) is assumed. Our most recent HST observation (2003 Feb. 2) showed a flux decrease of ∼ 20 percent from the optical peak. HST-1 is also flaring in x-rays, with our Chandra observations detecting a factor-of-2 increase in x-ray flux during mid-2002 and a decrease in later observations (Harris et al. 2003, Ap.J., in press;

http://xxx.lanl.gov/abs/astro-ph/?0302270). Optical monitoring of

M87 at sites with subarcsecond seeing (especially U-band monitoring to minimize background due to the host galaxy) is encouraged.”

2003 February 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green